An investment of ​$130 comma 000 was made by a business club. The investment was split into three parts and lasted for one year. The first part of the investment earned​ 8% interest, the second​ 6%, and the third​ 9%. Total interest from the investments was $ 9540. The interest from the first investment was 2 times the interest from the second. Find the amounts of the three parts of the investment.
The first part of the investment was ​$
nothing.

Answer by ankor@dixie-net.com(22740)

let the 3 investment amts = a, b, c
write an equation for each statement
:
An investment of ​$130,000 was made by a business club.
The investment was split into three parts and lasted for one year.
a + b + c = 130000
:
The first part of the investment earned​ 8% interest, the second​ 6%, and the third​ 9%. Total interest from the investments was $ 9540.
.08a + .06b + .09c = 9540
:
The interest from the first investment was 2 times the interest from the second.
.08a = 2(.06b)
.08a = .12b
a = .12b/.08
a = 1.5b
:
Find the amounts of the three parts of the investment.
substitute 1.5b for a in the first equation
1.5b + b + c = 130000
2.5b + c = 130000
c = 130000-2.5b
:
substitute .12b for .08a in the 2nd the equation
.12b + .06b + .09c = 9540
.18b + .09c = 9540
replace c with (130000-2.5b18)
.18b + .09(130000-2.5b) = 9540
.18b + 11700 - .225b = 9540
-.045b + 11700 = 9540
11700 - 9540 = .045b
2160 = .045b
b = 2160/.045
b = $48000 in b investment
then
c = 130000- 2.5(48000)
c = 130000 - 120000
c = $10000 in c investment
find a
130000 - 48000 - 10000 = $72000 in the a investment
:
:
:
Check the total interest
.08(72000) = 5760
.06(48000) = 2880
.09(10000) = 900
----------------------
total inter: $9540